In a school 74% like cricket, 76% like football, 82% like tennis. Then all the three sports is liked by at least how many students???

Grade:11

1 Answers

Askiitians_Expert Yagyadutt
13 years ago

Hii varsha !

 

It is based on a general formulae ...

 

n ( A U B U C ) = n ( A ) + n( B ) + n ( C ) - n(A int B ) - n ( A int C ) - n ( C int B )  + n ( A int B int C )

 

Now what we have to find out is....n ( A int B int C ) ...given  n(A) = 74  n(B) = 76  and n(C) = 82

 

Also n ( A U B U C ) = 300   (considering A B C = 100 % )

 

So  n ( A int B int C ) =   300 - ( 74 + 76 + 82 ) + n(A int B ) + n ( A int C ) + n ( C int B )

 

SO for case of atleast value....the unknown term can be taken as zero....

 

so  atleast value of n ( A int B int C )  = 300 - 232 = 68   ans

 

With reagrds

 

Yagya

askiitians_expert